if y1(t) and y2(t) are two solutions of the differential equation y^2-y'+y=0 then for any constants c1 and c2 c1y1(t)+c2y2(t) is also a solution

true or false and why
Answer #1

y (2) and y (+) are two solutions of the differential equation y2-yt y=0 then y, (+) satisfies y²-y+y=o y yś +42=0. y = dyMow %₂ = 84, + @Y 2 ² = (eg, + 94² y = 22 +43 z (@..)²+(94₂)²+28, 22 tagit ezeź) + egreb - 2)2+ winzeening @²m² + 3y + 296 Y